# On the utility of the switching theorem for adiabatic state preparation

### v1

@@ II Switching theorem
[…]
@@
While the switching theorem can estimate the true error in Eq. (1) very precisely, the error fluctuates a lot, which is the intrinsic property of the adiabatic theorem: rapid oscillations of phases cancel out excited states and only the ground state remains. In [25], it was suggested to consider the timescale-averaged quantity for errors. The “typical error” is:
[…]
@@
If $H^{m}(0)=H^{m}(1)=0$ for all $m\leq n$, the typical error from the switching theorem in the hyperadiabatic regime gives
$\bar{\epsilon}_{n}(T)=\frac{\bar{b}_{n}}{T^{n}}\;\;{\rm with}\;\;\bar{b}_{n}=\sqrt{\sum_{j\neq g}\left|\frac{\langle j(0)|H^{(n)}(0)|g(0)\rangle}{\Delta_{j,g}^{n+1}(0)}\right|^{2}+\sum_{j\neq g}\left|\frac{\langle j(1)|H^{(n)}(1)|g(1)\rangle}{\Delta_{j,g}^{n+1}(1)}\right|^{2}}\equiv\sqrt{(\bar{b}_{n}^{0})^{2}+(\bar{b}_{n}^{1})^{2}}.$ (4)
@@ III.1 Switching on of the switching theorem
For a given set of $\{b_{2}\,,\,b_{3}\,,\,\cdots\}$, this implies that the larger the value of $b_{1}$ (with the other $b_{n}$ fixed), the leading asymptotic behavior is realized for smaller values of $T$. This might naively suggest that increasing $b_{1}$ could be a way to observe the scaling behavior at smaller values of $T$. This is not sensible; increasing $b_{1}$ also increases the error, which is precisely the opposite of the goal.
@@ III.2 A Subtlety
where $k$ is a positive real number less than unity—and generally taken to be substantially less than unity. $g(s;k)$ has the properties that it has a vanishing first derivative with respect to $s$ and it approaches 1 when $s\gg k$. Thus, when $k\ll 1$, and $n$ is a positive integer $H_{k,n}(s)$ defined by
@@
has the property that it is very close to $H_{0}(s)$ for most of the region $10<s<$ 1, but it alters the behavior at the endpoints such that the first $n$ derivatives of $H$ with respect to $s$ vanish. Moreover, the lowest non-vanishing derivative at the endpoints for $H_{k,n}(s)$ is given by
@@
Therefore, in order for Eq. (9) to estimate the true typical error, $\bar{\epsilon}_{n}(T)\ll 1$, which corresponds to $T\gg(n!)^{1/2n}/\sqrt{k}$. It demonstrates that as the modified Hamiltonian Eq. (7) becomes close to the given Hamiltonian $H_{0}(s)$, i.e., $k\rightarrow 0$, it requires larger timescales to be in the hyperadiabatic regime. More importantly, as one modifies the Hamiltonian to have a better scaling behavior by increasing $n$, the timescale to be in the hyperadiabatic regime so that Eq. (9) can estimate the error precisely increases factorially.
@@ IV.1 Two-level system
Figure 2: Comparisons of errors, $\bar{\epsilon}_{T}$, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with $k=10^{-3}$ and $(E_{0},E_{1})=(1,1)$. The left plot shows the ratio of switching errors and the averaged true error with the solid line being $\bar{\epsilon}_{1}$ and the dotted line being $\bar{\epsilon}_{2}$. The right panel displays the three average errors with respect to the timescale. Note that the y-axis for the right plot is $\bar{\epsilon}$ times $T^{2}$, not just the error $\bar{\epsilon}$, for visualizing the scaling behavior.
@@
The model Hamiltonian that we study for adiabatic evolution may seem trivial in the sense that the initial and final states are the same. However, this feature is irrelevant for the questions that we want to answer, and the conclusions that follow would generally hold for non-trivial evolutions with different initial and final states, as is typical in realistic adiabatic state preparation setups. For presenting the plots, we will use $(E_{0},E_{1})=(1,1)$ such that $\Delta_{1,0}=1$ in arbitrary units at the endpoints. In this work, we are not concerned about the explicit dependence of the energy gaps in the Hamiltonian. The true error is obtained by solving the Schrödinger equation and calculated using Eq. (1) and averaged using Eq. (3), which will be denoted by the symbol $\bar{\epsilon}_{T}$. We will then compare it with the switching theorem estimate $\bar{\epsilon}_{n}(T)$ in Eq. (4).
@@
Here, $\frac{t}{t+k}\frac{1-t}{1-t+k}$ smoothens the behavior at the endpoints so that the first derivatives are zero when $k$ is not zero. The change in the intermediate path is small if $k$ is small. When $k$ is zero, the Hamiltonian is equal to $H_{0}(t)$, which, as discussed earlier, will show an error $\bar{\epsilon}\rightarrow\bar{b}_{1}/T$ as $T\rightarrow\infty$. The hope is that for sufficiently large but reasonable values of $T$, this minor modification will lead to a reduction in error. The $t$-independent multiplicative factor ensures that $f(1/2,k)=f_{0}(1/2)$, i.e., the intermediate behavior is left as close as possible to the original Hamiltonian.
Fig. 1 shows $f(t;k)$ for various $k$. Since it is even with respect to $t=0.5$, only its first half is plotted. It shows that when $k$ is small, $f(t;k)$ is nearly the same as $f_{0}(t)$ while its derivatives at the endpoints are 0. We want to stay as close as possible to the path traversed in the state space by $H_{0}(t)$ in the adiabatic limit, and therefore, $k$ is restrict to smaller values, which do not change the Hamiltonian path much.
@@
Figure 3: Comparisons of average errors, $\bar{\epsilon}_{T}$, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with two $k$, $10^{-4}$ and $10^{-3}$, and $(E_{0},E_{1})=(1,1)$. Note that switching errors can be larger than 1 since they are estimation, but true errors, Eq. (1), are always less than 1.
@@ IV.2 Three-level system
Figure 4: Comparisons of average errors, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with respect to $\bar{\epsilon}_{T}$, for the three-level system with identical off-diagonal elements at a fixed $k_{1}=k_{2}=k_{3}=10^{-4}$, while varying $T$.
Figure 5: Comparisons of the error ratio, $\bar{\epsilon}_{2}$/$\bar{\epsilon}_{T}$, by changing $k$ for the three-level system with identical off-diagonal elements, Eq. (16).
@@
In Fig. 5, the average error ratios $\bar{\epsilon}_{2}$/$\bar{\epsilon}_{T}$ are compared for different $k$. It shows that for large $k$, the average true errors follow $\bar{\epsilon}_{2}$ even at small $T$, whereas for small $k$, they gradually approach $\bar{\epsilon}_{2}$ as the timescale increases, similar to Fig. 3.
@@ V Extreme Scenario: Essential Singularity at the endpoints
with $\text{with }f_{\text{exp}}(t;k)$ =$=\begin{cases}t(1-t)\exp\left(-\frac{k}{t(1-t)}\right),&\text{if $0\leq t\leq 1$}\\ 0,&\text{otherwise}\end{cases}.$
[…]

### v3

# Practical limitations of the switching theorem for adiabatic state preparation

@@ II Switching theorem
where $\lVert\cdot\rVert$ denotes the $L^{2}$-norm and $|g_{i}\rangle$, $|g_{f}\rangle$ are the ground states of the initial and final Hamiltonians, respectively.
@@
While the switching theorem provides an asymptotically accurate estimate the true error in Eq. (1), the error fluctuates a lot, which is the intrinsic property of the adiabatic theorem: rapid oscillations of phases cancel out excited states and only the ground state remains. In [25], it was suggested to consider the timescale-averaged quantity for errors. The “typical error” is:
$\bar{\epsilon}(T)\equiv\frac{1}{2\sqrt{T\tau_{0}}}\int_{T-\sqrt{T\tau_{0}}}^{T+\sqrt{T\tau_{0}}}dT^{\prime}\,\epsilon(T^{\prime})$ (3)
@@
If $H^{m}(0)=H^{m}(1)=0$ for all $m\leq n$, Ref. [25] showed that the typical error from the switching theorem in the hyperadiabatic regime gives
$\bar{\epsilon}_{n}(T)=\frac{\bar{b}_{n}}{T^{n}}\;\;{\rm with}\;\;\bar{b}_{n}=\sqrt{\sum_{j\neq g}\left|\frac{\langle j(0)|H^{(n)}(0)|g(0)\rangle}{\Delta_{j,g}^{n+1}(0)}\right|^{2}+\sum_{j\neq g}\left|\frac{\langle j(1)|H^{(n)}(1)|g(1)\rangle}{\Delta_{j,g}^{n+1}(1)}\right|^{2}}\equiv\sqrt{(\bar{b}_{n}^{0})^{2}+(\bar{b}_{n}^{1})^{2}},$ (4)
where the superscript $n$ on $H$ denotes the $n$th derivative of the Hamiltonian, and $|g(t)\rangle$ $|j(t)\rangle$ denotes the instantaneous ground state and the $j$th excited state, respectively. Additionally, $\Delta_{j,g}$ is the energy gap between the $j$th excited state and the ground state, i.e., $\Delta_{j,g}(t)=E_{j}(t)-E_{g}(t)$.
Moreover, in the hyperadiabatic regime, the typical error has the following inequality:
$\epsilon(T)\leq\sqrt{2}\bar{\epsilon}(T),$ (5)
which was shown in [25]. Therefore, the hyperadiabatic regime is important in the sense that the error has a definite scaling behavior in $T$ and it can be bounded from the typical error.
@@ III.1 Switching on of the switching theorem
For a given set of $\{b_{2}\,,\,b_{3}\,$$\cdots\},$ this implies that the larger the value of $b_{1}$ (with the other $b_{n}$ fixed), the leading asymptotic behavior is realized for smaller values of $T$. This might naively suggest that increasing $b_{1}$ could be a way to observe the scaling behavior at smaller values of $T$. This is not sensible; increasing $b_{1}$ also increases the error, which is precisely the opposite of the goal.
@@ III.2 A subtlety
where $k$ is a positive real number less than unity—and generally taken to be substantially less than unity. $g(s;k)$ has the properties that it is zero when $s=0$ and it approaches 1 when $s\gg k$. Thus, when $k\ll 1$, and $n$ is a positive integer $H_{k,n}(s)$ defined by
@@
has the property that it is very close to $H_{0}(s)$ for most of the region $0<s<1$, but it alters the behavior at the endpoints such that the first $n$ derivatives of $H$ with respect to $s$ vanish. Moreover, the lowest non-vanishing derivative at the endpoints for $H_{k,n}(s)$ is given by
@@
Therefore, for a given $n$, in order for Eq. (10) to estimate the true typical error, the evolution should be in the hyperadiabatic regime, i.e., $\bar{\epsilon}_{n}(T)\ll 1$, which corresponds to $T\gg(n!)^{1/2n}/\sqrt{k}$. It demonstrates that as the modified Hamiltonian Eq. (8) becomes close to the given Hamiltonian $H_{0}(s)$, i.e., $k\rightarrow 0$, it requires larger timescales to be in the hyperadiabatic regime. More importantly, as one modifies the Hamiltonian to have a better scaling behavior by increasing $n$, the timescale to be in the hyperadiabatic regime so that Eq. (10) can estimate the error precisely increases factorially.
@@ IV.1 Two-level system
Figure 2: Comparison of errors, $\bar{\epsilon}_{T}$, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with $k=10^{-3}$ and $(E_{0},E_{1})=(1,1)$. The left plot shows the ratio of switching errors and the averaged true error with the solid line being $\bar{\epsilon}_{1}$ and the dotted line being $\bar{\epsilon}_{2}$. The right panel displays the three average errors with respect to the timescale. Note that the y-axis for the right plot is $\bar{\epsilon}$ times $T^{2}$, not just the error $\bar{\epsilon}$, for visualizing the scaling behavior.
@@
The model Hamiltonian that we study for adiabatic evolution may seem trivial in the sense that the initial and final states are the same. However, this feature is irrelevant for the questions that we want to answer, and the conclusions that follow would generally hold for non-trivial evolutions with different initial and final states, as is typical in realistic adiabatic state preparation setups. For presenting the plots, we will use $(E_{0},E_{1})=(1,1)$ such that $\Delta_{1,0}=1$ in arbitrary units at the endpoints. In this work, we are not concerned about the explicit dependence of the energy gaps in the Hamiltonian [^1] The focus of this work is on the relative delay in the onset of the asymptotic caused by endpoint modification, while the energy gaps are fixed at the endpoint.. The true error is obtained by solving the Schrödinger equation [^2] The time-dependent Schrödinger equation is solved using LSODA algorithm implemented on Mathematica. and calculated using Eq. (1) and averaged using Eq. (3), which will be denoted by the symbol $\bar{\epsilon}_{T}$. We will then compare it with the switching theorem estimate $\bar{\epsilon}_{n}(T)$ in Eq. (4).
@@
Here, $\frac{t}{t+k}\frac{1-t}{1-t+k}$ smoothens the behavior at the endpoints so that the first derivatives are zero when $k$ is not zero. The change in the intermediate path is small if $k$ is small. When $k$ is zero, the Hamiltonian is equal to $H_{0}(t)$, which, as discussed earlier, will show an error $\bar{\epsilon}\rightarrow\bar{b}_{1}/T$ as $T\rightarrow\infty$. The objective is to determine whether, for sufficiently large but reasonable values of $T$, this minor modification will lead to a reduction in error. The $t$-independent multiplicative factor ensures that $f(1/2,k)=f_{0}(1/2)$, i.e., the intermediate behavior is left as close as possible to the original Hamiltonian.
Fig. 1 shows $f(t;k)$ for various $k$. Since it is even with respect to $t=0.5$, only its first half is plotted. It shows that when $k$ is small, $f(t;k)$ is nearly the same as $f_{0}(t)$ while its derivatives at the endpoints are 0. We want to stay as close as possible to the path traversed in the state space by $H_{0}(t)$ in the adiabatic limit, and therefore, $k$ is restricted to smaller values, which do not change the Hamiltonian path much.
@@
Figure 3: Comparison of average errors, $\bar{\epsilon}_{T}$, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with two $k$, $10^{-4}$ and $10^{-3}$, and $(E_{0},E_{1})=(1,1)$. Note that switching errors can be larger than 1 since they are estimation, but true errors, Eq. (1), are always less than 1.
@@ IV.2 Three-level system
Figure 4: Comparison of average errors, $\bar{\epsilon}_{1}$, and $\bar{\epsilon}_{2}$, with respect to $\bar{\epsilon}_{T}$, for the three-level system with identical off-diagonal elements at a fixed $k_{1}=k_{2}=k_{3}=10^{-4}$, while varying $T$.
Figure 5: Comparison of the error ratio, $\bar{\epsilon}_{2}$$\bar{\epsilon}_{T}$, by changing $k$ for the three-level system with identical off-diagonal elements, Eq. (17).
@@
In Fig. 5, the average error ratios $\bar{\epsilon}_{2}$$\bar{\epsilon}_{T}$ are compared for different $k$. It shows that for large $k$, the average true errors follow $\bar{\epsilon}_{2}$ even at small $T$, whereas for small $k$, they gradually approach $\bar{\epsilon}_{2}$ as the timescale increases, similar to Fig. 3.
@@ V Extreme Scenario: Essential Singularity at the endpoints
$\text{with }f_{\text{exp}}(t;k)$ $=\begin{cases}t(1-t)\exp\left(-\frac{k}{t(1-t)}\right),&\text{if $0\leq t\leq 1$}\\ 0,&\text{otherwise}\end{cases}.$
As $k$ approaches zero, the difference between $H(t,k)$ in Eq. (19) and $H_{0}(t)$ in Eq. (11) at all intermediate times diminishes. The switching theorem is not particularly useful in estimating the error for the Hamiltonian in Eq. (19) when $k\neq 0$ since its derivatives are zero at the endpoints. One can always calculate the derivatives of $H$ to see that all the derivatives become zero at the endpoints. We cannot use the switching theorem here to make estimates for the errors. However, we can draw inspiration from the switching theorem to motivate such a modification to the Hamiltonian at its endpoints. This can be done by viewing the $f(t,k)$ as a limit of analytic functions with well-defined derivatives at the endpoints. The function $f_{\rm exp}(t,k)$ defined in Eq. (19) is compared against $f_{0}(t)$ defined in Eq. (12) in Fig. 1. We compare the final error after time evolution with $H(t)$ from Eq. (19) with $k=10^{-4}\,,\,10^{-3}\,,\,10^{-2}$ to the error accumulated during the Hamiltonian evolution governed by Eq. (11).
@@ VI Discussion
Although the numerical demonstrations in this work utilized low-dimensional models, the fundamental trade-off identified is expected to be universal for gapped quantum systems. The switching theorem derives its error estimates primarily from the behavior of the Hamiltonian at the endpoints, where ground and low-lying states contribute most significantly, rather than the specific dimensionality of the Hilbert space. Consequently, the ‘practical limitations’ observed in our simple models are not artifacts of low dimensionality but are intrinsic features of the switching theorem, implying that more complex, realistic many-body Hamiltonians will confront the same barriers to efficient error reduction.
@@
Author contributions
All authors contributed significantly and in complementary ways to this work. AL conducted the initial study of the systems as part of a high school summer project under the guidance of HO, MP, and TC. HO drafted the initial version of the manuscript, generated the figures, and participated in manuscript revisions. MP supervised the project, contributed to writing, and assisted in editing. TC oversaw the project throughout all stages. The contributions were balanced, with each author playing a key role in the completion of the study.
Data availability statement
Data will be made available on reasonable request.